Find two consecutive even integers such that the square of the second, decreased by twice the first, is 52.
Answer by AnlytcPhil(1810)

Find two consecutive even integers such that the square of the second,
decreased by twice the first, is 52.
Let the smaller even integer be N. Then the next consecutive even
integer to N is N+2
>>...the square of the second, decreased by twice the first, is 52...<<
Translation: (N + 2)2 - 2N = 52
(N + 2)2 - 2N = 52
N2 + 4N + 4 = 2N + 52
N2 + 2N - 48 = 0
(N + 8)(N - 6) = 0
N + 8 = 0, so N = -8
N - 6 = 0, so N = 6
So the integers, N and N+2, are either -8 amd -6, or 6 and 8
Edwin